Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

Wilkinstown Open 2

Distribution

Dr. Trevor West DUCAC Chairman
Aubrey Storey DUCAC Secretary
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record
Susan Thompson Trinity News

Summary

Venue Wilkinstown Target Shooting Club, Wilkinstown
Date March 9-10, 2002
Discipline 10m Air Rifle

DURC Attendees

Mark Dennehy Kate Fanning
Ray Reilly  Keith White

Scores

Mark Dennehy 555
Ray Reilly 547
Kate Fanning 383

Prizes Won

Mark Dennehy 6th place in Finals

Report

Despite inclement weather and horrible driving conditions, DURC sent a small team to the WTSC Open this weekend. Shooting conditions were less than optimal as the weather caused the light in the range to vary from bright to dull, which of course upset the sight picture for the shooters. Kate Fanning was particularly offput by this as she has not shot in outdoor .22 competitions yet, which is the only place where it is possible to learn to compensate for such problems.

Mark Dennehy managed to compensate well, and set a new personal best of 555 points, and placed 6th in the final, fighting off Helen Meade for the second competition in a row.

Ray Reilly is also showing progress back to his original form with a 541. Ray was forced to retire from air shooting for almost a year with a back injury, but hopefully will now make a comeback to the scene.

And finally, Keith White’s software for administering competitions, developed soley by him, is nearing it’s first release. Allready it has become the standard software for this purpose in Ireland, and he should be congradulated on this as he has put nearly 8 years of development into it, with extensive user testing.

The next competition in Air Rifle in Ireland is the Nationals - lots of challanges have been issued for this! First however, there is the Isle of Man away trip at the end of this month.

Misé le meas,

Mark Dennehy, DURC Secretary